OpenStudy (anonymous):

Determine all integer solutions to 5x - 7y = 99. Professor says to watch the minus sign, having trouble with that.

OpenStudy (asnaseer):

this is very similar again to the ones you've already done

OpenStudy (asnaseer):

where are you getting stuck?

OpenStudy (anonymous):

I'm getting stuck on the negative sign. AT what time during the continued fraction do I address the minus sign.

OpenStudy (asnaseer):

continued fraction? I just use simple algebra.

OpenStudy (asnaseer):

look at what I did in your question here: http://openstudy.com/updates/4fe615bee4b02c91101af1b3

OpenStudy (anonymous):

Professor's example for x+y=z formulas for finding all integer solutions shows continued fraction, then using the solution obtained to plug into a formula for finding all (n) solutions.

OpenStudy (asnaseer):

do you have a web site link that describes this method?

OpenStudy (anonymous):

it's in our book. I also came up with a different answer that seems to work for all (n) for the other problem, I'll post it there.

OpenStudy (asnaseer):

I haven't heard of the continued fraction method which is why I was interested in learning it

OpenStudy (anonymous):

using 2x+11y=10

OpenStudy (asnaseer):

but I have solved this using the algebraic method I used in the previous question

OpenStudy (anonymous):

what would the algebreic method yeild, and how would you do it?

OpenStudy (asnaseer):

shall I show you the solution to 2x+11y=10 using algebraic method?

OpenStudy (anonymous):

continued fraction first asks for gcd (a,b), continued fraction table looks like this 77 14 7 0 5 2 0 1 5 11 1 0 1 12

OpenStudy (asnaseer):

where did the 77, 14, and 7 come from in the 1st row?

OpenStudy (asnaseer):

ah! King George is here!

OpenStudy (asnaseer):

King George - do you know about this "continued fraction" method - and where can I learn more about it?

OpenStudy (anonymous):

i'd rather learn what to do in case of minus sign for continued fraction, which is what prof wants

OpenStudy (kinggeorge):

I've never seen the continued fraction method.

OpenStudy (asnaseer):

@tim93 is there a formal name for this method?

OpenStudy (anonymous):

77 = a, 14 = b, 7 is the remainder when 14 goes into 77, it continues out until 0 is reached

OpenStudy (anonymous):

the next line represents the integer quotient

OpenStudy (anonymous):

diophantine equations, euclidian algorithm for gcd, discrete math course

OpenStudy (anonymous):

The bottom two rows are generated starting with 0,1 on first row, 1,0 on last row

OpenStudy (asnaseer):

ok - let me try and find more information on this method and then I'll try and see if I can solve this using this new method.

OpenStudy (kinggeorge):

Oh wait, I have seen this. I've heard it called "The box method" or "The magic box method"

OpenStudy (anonymous):

to find the next values, multiply the number above by the number to the left, and add the number two to the left

OpenStudy (asnaseer):

thanks @KingGeorge that might help me find more information on this.

OpenStudy (asnaseer):

@tim93 do you want to see how I solved this using the algebraic method as well or shall I come back once I've learnt this new continued fraction method?

OpenStudy (anonymous):

now, start with positive on the bottom left, and count alternating negative and positive until you reach the second to last value on the right, this value corresponds to the upper left value, so you're left with one solution to the problem of 77*10 + 14*-50 = 70 (multiplied by ten because we had calculated one solution for 7 in the box)

OpenStudy (anonymous):

yes, i'd love to see the algebraic solution, but for the original questuion in this thread, as you answered the other on the other thread, btw, thanks for that

OpenStudy (asnaseer):

ok, here goes...

OpenStudy (asnaseer):

\[\begin{align} 5x-7y&=99\\ \therefore x&=\frac{7y+99}{5}=\frac{2y+5y+4+95}{5}=\frac{2y+4}{5}+y+19\\ \implies 2y+4&=5n\\ \therefore y&=\frac{5n-4}{2}\implies n\text{ must be even}\\ \therefore n&=2m\\ \therefore y&=\frac{10m-4}{2}=5m-2 \end{align}\]

OpenStudy (asnaseer):

that gives all integer y's

OpenStudy (asnaseer):

now just substitute that back into the expression for x

OpenStudy (asnaseer):

to get:\[x=7m+17\]

OpenStudy (asnaseer):

so final solution is: x = 7m + 17 y = 5m - 2 for all integer m

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

how did you get from second to third line, and what happens to y + 19

OpenStudy (asnaseer):

ok, you understand this part:\[x=\frac{2y+4}{5}+y+19\]

OpenStudy (anonymous):

solved for x, and factored

OpenStudy (asnaseer):

for x to be an integer, the fractional component on the RHS must also be an integer - agreed?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

therefore (2y+4) must be a multiple of 5

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

that leads to: 2y + 4 = 5n

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

clearer

OpenStudy (asnaseer):

all clear now?

OpenStudy (anonymous):

in essence, it's inductive in reasoning? How to make the leap from the original equation to the second and third lines takes some practice i guess

OpenStudy (asnaseer):

yes - you are right - with lots of practice you learn to recognise things like this more easily

OpenStudy (anonymous):

could we have left it without saying that n must be even?

OpenStudy (asnaseer):

I don't think so, because we end up with:\[y=\frac{5n-4}{2}\]which is not an integer for all n

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It will come in handy for me too, as this guy explains things somewhat briefly, and we are expected to learn beyond the six pages per chapter

OpenStudy (kinggeorge):

I don't know if it says in that link or not, but the magic box method only works when the gcd is 1.

OpenStudy (asnaseer):

so the factors of x and y have to be coprime?

OpenStudy (anonymous):

but it seemed to work for the equation in the other problem

OpenStudy (kinggeorge):

The factors must be coprime.

OpenStudy (asnaseer):

ok - well I have plenty of reading to do then thanks to you two! :D I'll get back to you once I have learnt this method.

OpenStudy (kinggeorge):

Or at least, that's what I was taught in my classes.

OpenStudy (anonymous):

once the cofactors are coprime, then once the box is solved, we'd multiply by some x to revert back to original equation maybe?

OpenStudy (kinggeorge):

What was the other equation?

OpenStudy (anonymous):

77 14 7 0 is the first line of the box, with 7 being gcd(77,14)

OpenStudy (anonymous):

oops, i guess gcd has nothing to do with that line, it's just the remainder

OpenStudy (anonymous):

coincidentally gcd

OpenStudy (anonymous):

i would love it if someone showed me the box method for a negative as one of the factors, once a solution is reached, our prof gave us these two formulas, where s and t are one solution for x and y respectively, and d is gcd(a,b), and k is any integer x=s + (a/d)k y=t-(b/d)k

OpenStudy (kinggeorge):

I'll work some stuff out and see if I can get a working example with some negative numbers. First though, what exactly do you mean by "a negative as one of the factors?" Do you mean something like \(23x-32y=1\)?

OpenStudy (anonymous):

yes, and how does it fit into the box, if at all, perhaps it is disregarded until the next step

OpenStudy (kinggeorge):

I've got to leave for a bit, but I'll post some stuff on that when I get back.

OpenStudy (anonymous):

thanks George, for your dedication to MATH

OpenStudy (kinggeorge):

I'll just put an example for \(23x+32y=1\) here. First, Euclidean algorithm.\[32=23\cdot1+9\]\[23=9\cdot2+5\]\[9=5\cdot1+4\]\[5=4\cdot1+1\]\[4=1\cdot4+0\]So the gcd is 1. Now create the table \[\begin{array}{|c|c|c|c|c|c|c|} \hline &&1&2&1&1&4\\ \hline 0&1&1&2+1=3&3+1=4&4+3=7&28+4=32 \\ \hline 1&0&1&2&2+1=3&3+2=5&20+3=23 \\ \hline \end{array}\]Therefore, you know that a solution is \[7(23)-5(32)=1.\]Notice that in the last two columns, we are multiplying diagonally to achieve the solution. Diagonally down will be positive, diagonally up will be negative.

OpenStudy (anonymous):

Wouldn't you have to start the Euclidian algorith with 7 and -5 for the original equation, not 23 and 32?

OpenStudy (kinggeorge):

No, we use the Euclidean Algorithm to find the gcd of 32, 23, and using the successive quotients, we can compute a solution to \(23x+32y=1\).

OpenStudy (anonymous):

how would you draw up 5x - 7y = 99

OpenStudy (kinggeorge):

Well, let's see here. \[7=5\cdot1+2\]\[5=2\cdot2+1\]\[2=1\cdot2+0\]Draw the table.\[\begin{array}{|c|c|c|c|c|} \hline &&1&2&2\\ \hline 0&1&1&2+1=3&6+1=7\\ \hline 1&0&1&2&4+1=5\\ \hline \end{array}\]So \(3(5)-2(7)=1\).

OpenStudy (kinggeorge):

To find \(5x-7y=99\), just multiply by 99. That means \[5(297)-7(198)=99\]

OpenStudy (anonymous):

did you use the magic box?

OpenStudy (kinggeorge):

That may not be the simplest solution, but it is a solution. I used the magic box to find a solution to \(5x-7y=1\). With that solution, I found the solution to \(5x-7y=99\).

OpenStudy (anonymous):

so, in the box, do you use 7 or -7

OpenStudy (kinggeorge):

I'm using 7 in the box.

OpenStudy (kinggeorge):

You just have to remember that the very last multiplication, which is diagonally up, you should multiply by -1.